# A Gentle Introduction to the Axiom of Choice

### v2

@@ 1 Introduction
Described by David Hilbert as the axiom “most attacked up to the present in the mathematical literature” [Hil26], the axiom of choice ($\mathsf{AC}$) indeed has a fascinating history (see [Moo82]). In 1883, Georg Cantor had proposed the well-ordering principle, i.e. that every set can be well-ordered, [^1] We say that a set $X$ can be well-ordered if there is a well-ordering on $X$, i.e. a binary relation $\preceq$ on $X$ that is reflexive ($\forall x\in X,x\preceq x$), transitive ($\forall x,y,z\in X$, if $x\preceq y$ and $y\preceq z$, then $x\preceq z$), anti-symmetric ($\forall x,y\in X$, if $x\preceq y$ and $y\preceq x$, then $x=y$), total ($\forall x,y\in X$, $x\preceq y$ or $y\preceq x$), and well-founded (every nonempty subset of $X$ has a $\preceq$-minimal element). For example, the standard ordering $\leqslant$ on the set $\mathbb{N}=\{0,1,2,\ldots\}$ of natural numbers is a well-ordering, and the standard ordering $\leqslant$ on the set $\mathbb{Z}=\{\ldots,-2,-1,0,1,2,\ldots\}$ of integers fails only to satisfy the last condition, since, for example, $\mathbb{Z}$ itself has no $\leqslant$-minimal element. (Note, however, that the integers can be well-ordered by, for example, constructing your favorite bijection of sets $h:\mathbb{N}\to\mathbb{Z}$ and ordering the integers by the induced ordering from $h$ and the standard ordering $\leqslant$ on $\mathbb{N}$.) Similarly, the standard ordering $\leqslant$ on the set $0,\infty)$ of non-negative real numbers—which itself has a $\leqslant$-minimal element—fails to satisfy the last condition since, for example, the subset $(0,\infty)$ of positive real numbers has no $\leqslant$-minimal element. (Unfortunately, one must work harder to prove that there is a well-ordering on $0,\infty)$. Indeed, the initial motivation behind Cantor’s well-ordering principle came from the question of whether the set $\mathbb{R}$ of real numbers can be well-ordered—discussed in detail in [Moo82].) without proof, as a valid law of thought. Since his so-called law was not immediately accepted, Cantor found himself seeking a proof. It was in order to provide such a proof that $\mathsf{AC}$ was first explicitly formulated by Ernst Zermelo in 1904 [Zer04].
@@
It is also worth explicitly noting that, even if an argument (in this case, an existence proof) seems to rely on making arbitrary “choices,” $\mathsf{AC}$ may not be needed. For instance, a standard proof of the Bolzano-Weierstrass theorem—every bounded sequence $(a_{i})_{i\in\mathbb{N}}$ in $\mathbb{R}$ contains a convergent subsequence—relies on iteratively “choosing” infinitely many terms $(a_{i_{j}})_{j\in J\subseteq\mathbb{N}}$ from $(a_{i})_{i\in\mathbb{N}}$ satisfying some properties (see [Abb15, Theorem 2.5.5]). In this case, since the indexing set $\mathbb{N}$ is well-ordered by $\leqslant$ (see footnote 1), each “choice” can be made definably by choosing the lowest-index term of the sequence satisfying those properties.
@@
We use the abbreviation $\mathsf{ZFC}$ for Zermelo-Fraenkel set theory with $\mathsf{AC}$—which is the standard axiomatic basis for mathematics—and the abbreviation $\mathsf{ZF}$ for Zermelo-Fraenkel set theory without assuming $\mathsf{AC}$. However, familiarity with the axioms of $\mathsf{ZF}$ is not assumed or necessary for this discussion.
@@ 2.2 Mathematical inconveniences
Often, a reason to not believe $\mathsf{AC}$ is its seemingly paradoxical consequences. If there were such a thing as a canonical [^4] See, for example, “the most spectacular [counterintuitive geometrical consequence of $\mathsf{AC}$]” in [Bel09, Ch.I] and “the most stunning [paradoxical result that is demonstrable in $\mathsf{ZFC}$ but not in $\mathsf{ZF}$]” in [Her06, §5.2], as well as [JW96, §9.4: Banach-Tarski paradox] and [Jec73, §1.3: A paradoxical decomposition of the sphere]. example of a “disastrous” consequence of the axiom of choice, it would be the Banach-Tarski paradox—which, roughly stated, says that a closed three-dimensional ball $B$ can be decomposed into finitely many pieces, which can be rearranged into two disjoint copies of $B$ using only rigid transformations.
@@
Banach-Tarski paradox ($\mathsf{BT}$). Assuming $\mathsf{AC}$, for a closed three-dimensional ball $B$, there exists a decomposition $B=B_{1}\sqcup B_{2}$ such that $B_{1}\approx B\approx B_{2}$ [BT24]. [^5] For a detailed discussion on $\mathsf{BT}$, see [TW16].
@@
$\mathsf{BT}$ (hence also $\mathsf{AC}$) implies another mathematically inconvenient result: the existence of subsets of $\mathbb{R}^{n}$ that are not (Lebesgue) measurable. [^6] For a formal and self-contained treatment of the Lebesgue measure on $\mathbb{R}^{n}$, which is beyond the scope of this article, see [SS05, Ch.1]. In the case of $\mathbb{R}$, the intuition behind (Lebesgue) measurability comes from the ordinary notion of length: the measure of a point is zero and that of an open interval $(a,b)\subseteq\mathbb{R}$ is its length, denoted m(a,b)=bam(a,b)=b-a. Since any open subset $U$ of $\mathbb{R}$ can be written uniquely as a disjoint union of open intervals [SS05, Theorem 1.3], the measure of $U$ can be defined as the sum of the lengths of these open intervals. Moreover, the Borel sets—defined as the smallest collection of subsets of $\mathbb{R}$ that includes the open sets and is closed under taking complements and countable unions—are measurable [SS05, Ch.1, §3]. (For instance, since singletons are closed, hence Borel, all countable sets are Borel, hence measurable.) A subset $A$ of $\mathbb{R}$ has measure zero if it can be covered by arbitrarily small unions of open intervals: for any $\varepsilon>0$, there are open sets $U_{i}$ with $\sum_{i}m(U_{i})<\varepsilon$ such that $A\subseteq\bigcup_{i}U_{i}$. (For instance, countably infinite sets have measure zero.) More generally, a set $A\subseteq\mathbb{R}$ is measurable if and only if it it differs from a Borel set by a set of measure zero [SS05, Corollary 3.5]. More specifically, as a consequence of $\mathsf{BT}$, there cannot exist a finitely additive measure that is defined on all subsets of $\mathbb{R}^{3}$ and invariant under rigid transformations.
@@
The standard example of a non-measurable subset of $\mathbb{R}$ can be defined as follows. For $x,y\in\mathbb{R}$, we write $x\sim_{\mathbb{Q}}y$ to mean that $x-y\in\mathbb{Q}$, where $\mathbb{Q}$ denotes the set of rational numbers. This is an equivalence relation on $\mathbb{R}$. [^7] A binary relation $\sim$ on a set $X$ is said to be an equivalence relation if it is reflexive (see footnote 1), symmetric ($\forall x,y\in X$, if $x\sim y$ then $y\sim x$), and transitive (again, see footnote 1). Given an equivalence relation $\sim$ on $X$, the $\sim$-equivalence class of an element $x\in X$ is the set $x]_{\sim}=\{y\in X\mid x\sim y\}$, and the partition of $X$ by $\sim$ is the set $X/{\sim}=\{[x]_{\sim}\mid x\in X\}$ of all $\sim$-equivalence classes. Since each $\sim$-equivalence class is nonempty (as $x\in[x]_{\sim}$), $\mathsf{AC}$ guarantees the existence of a choice function on the partition $X/{\sim}$. We define a choice set on $X/{\sim}$ as the range of such a choice function. Since distinct equivalence classes are disjoint, a choice set on $X/{\sim}$ will contain exactly one element from each equivalence class. A Vitali set $V$ is a subset of $[0,1]$ containing exactly one element from each $\sim_{\mathbb{Q}}$-equivalence class, or simply a choice set on $0,1]/{\sim_{\mathbb{Q}}}$ (see footnote 7), whose existence is guaranteed by $\mathsf{AC}$. A standard argument of the non-measurability of such a $V$, relying on the translation-invariance of measure, can be found in [SS05, Ch.1, §3].
@@
Note, however, that the existence of a non-measurable set is not provable in $\mathsf{ZF}$, i.e. it is consistent with $\mathsf{ZF}$ (assuming the consistency of inaccessible cardinals) that all subsets of $\mathbb{R}$ are measurable [Sol70]. [^8] For a brief discussion on assumptions that are consistent with $\mathsf{ZF}$ and incompatible with $\mathsf{AC}$, in the context of analysis, see [Sch97, §14]. In this regard, it would seem as though the natural way of eliminating $\mathsf{BT}$ would be to insist that all subsets of $\mathbb{R}$ (hence all subsets of $\mathbb{R}^{n}$ [SS05, Proposition 3.6]) are measurable and to suitably weaken $\mathsf{AC}$. However, this assumption leads to another paradox, which we describe briefly, that is just as unsettling as the one being eliminated.
@@ 3.1 Because a lot of math follows from $\mathsf{AC}$ or needs $\mathsf{AC}$
Thus far, we have seen some uses of $\mathsf{AC}$ that are not necessarily intuitive. For instance, in the introduction, we mentioned that $\mathsf{AC}$ was used to prove the well-ordering principle. In fact, this use is necessary, in the sense that $\mathsf{AC}$ equivalent to the well-ordering principle. $\mathsf{AC}$ is also regularly used throughout mathematics in another equivalent form, namely Zorn’s lemma: if $(P,\preceq)$ is a nonempty partially ordered set in which every nonempty chain $C\subseteq P$ has a $\preceq$-upper bound, then $P$ has a $\preceq$-maximal element. [^11] A binary relation $\preceq$ on $P$ is said to be a partial order if it is reflexive, anti-symmetric, and transitive (see footnote 1). A subset $C$ of a partially ordered set $(P,\preceq)$ is said to be a chain if the restriction of $\preceq$ to $C$ is total (again, see footnote 1). The equivalence of these statements may not seem obvious at all, as reflected by a famous (ironic) quote of Jerry Bona from 1977: “The Axiom of Choice is obviously true; the Well-Ordering principle is obviously false; and who can tell about Zorn’s Lemma?” [Sch97]. The curious reader, who at this point has surely attempted showing that these statements are equivalent, may find a proof in [Jec73, Theorem 2.1].
@@
Assuming $\mathsf{AC}$, these notions agree: $g$ is continuous at $x$ if and only if $g$ is sequentially continuous at $x$. However, the backward direction of this equivalence cannot be proved in $\mathsf{ZF}$, i.e. it is consistent with $\mathsf{ZF}$ that the backwards direction fails ([Jae65]; discussed in [Moo82, Ch.1.2]).
A natural question to raise is whether we need the “full strength” of $\mathsf{AC}$ for this equivalence to hold. This turns out not to be the case: the backwards direction is equivalent to countable choice from sets of reals, i.e. the assertion that any countable family of nonempty subsets of $\mathbb{R}$ admits a choice function, abbreviated $\mathsf{CC}(\mathbb{R})$ [Her06, Theorem 4.54, 8 and 10]. $\mathsf{CC}(\mathbb{R})$ is the restriction of $\mathsf{AC}$ where $\mathcal{F}$ is countable, i.e. $\lvert\mathcal{F}\rvert\leqslant\lvert\mathbb{N}\rvert$, and each element of $\mathcal{F}$ is a nonempty subset of $\mathbb{R}$. This is an obvious consequence of $\mathsf{AC}$ restricted to countable families of arbitrary nonempty sets, abbreviated $\mathsf{CC}$, which itself is strictly weaker than $\mathsf{AC}$. [^12] Both of the implications $\mathsf{AC}\implies\mathsf{CC}\implies\mathsf{CC}(\mathbb{R})$ are strict. See [Jec73, Ch.8] for the irreversibility of the first implication. Since $\mathsf{CC}(\mathbb{R})$ holds in every permutation model and is a boundable statement (see [HR98, III.§2] and [Jec73, Ch.6])—hence, transferrable to a $\mathsf{ZF}$ model—the irreversibility of the second implication can be seen through any permutation model where $\mathsf{CC}$ fails, such as the Basic Fraenkel Model (see [HR98, III.§2.1] or [Jec73, Ch.4.3]).
Note, furthermore, that $\mathsf{CC}(\mathbb{R})$ implies that the Lebesgue measure (see footnote 6) is countably additive, i.e. if $(A_{i})_{i\in\mathbb{N}}$ are pairwise disjoint measurable subsets of $\mathbb{R}$, then $m(\bigcup_{i}A_{i})=\sum_{i}m(A_{i})$ [Her06, §5.1(E13)]. Countable additivity is not provable in $\mathsf{ZF}$, i.e. it is consistent with $\mathsf{ZF}$ that the Lebesgue measure is not countably additive. [^13] For example, in the Feferman-Lévy Model ([FL63]; see [HR98, III.§1.9]), $\mathbb{R}$ is a countable union of countable sets, so there is no nontrivial countably additive measure on $\mathbb{R}$. (Of course, this in turn means that $\mathsf{CC}(\mathbb{R})$ fails in this model, and so does the aforementioned equivalence of continuity notions.)
@@
Let $F$ be a field and $V$ be an $F$-vector space. If $V$ has a finite spanning set, i.e. if there are $x_{1},\ldots,x_{k}\in V$ such that any $y\in V$ can be written as a linear combination $y=\sum_{i=1}^{k}c_{i}x_{i}$, with $c_{i}\in F$, then it is not hard to show, without appealing to $\mathsf{AC}$, that $V$ has a (finite) basis. Furthermore, if an $F$-vector space $V$ has a finite basis, it is once again not hard to show, without appealing to $\mathsf{AC}$, that any two bases of $V$ must have the same cardinality—so the dimension of $V$ can be defined as the cardinality of any basis of $V$.
@@
For results and discussions on results in other fields of mathematics that rely on $\mathsf{AC}$ or its consequences, we refer the interested reader to [Jec73, Ch.10], [Hod74], [Moo82], [GK91], [HR98] and references there, and [Her06, §4].

### v3

@@ 1 Introduction
Described by David Hilbert as the axiom “most attacked up to the present in the mathematical literature” [Hil26], the axiom of choice ($\mathsf{AC}$) indeed has a fascinating history (see [Moo82]). In 1883, Georg Cantor had proposed the well-ordering principle, i.e., that every set can be well-ordered [^1] We say that a set $X$ can be well-ordered if there is a well-ordering on $X$, i.e., a binary relation $\preceq$ on $X$ that is reflexive ($\forall x\in X,x\preceq x$), transitive ($\forall x,y,z\in X$, if $x\preceq y$ and $y\preceq z$, then $x\preceq z$), anti-symmetric ($\forall x,y\in X$, if $x\preceq y$ and $y\preceq x$, then $x=y$), total ($\forall x,y\in X$, $x\preceq y$ or $y\preceq x$), and well-founded (every nonempty subset of $X$ has a $\preceq$-minimal element). For example, the standard ordering $\leqslant$ on the set $\mathbb{N}=\{0,1,2,\ldots\}$ of natural numbers is a well-ordering, and the standard ordering $\leqslant$ on the set $\mathbb{Z}=\{\ldots,-2,-1,0,1,2,\ldots\}$ of integers fails only to satisfy the last condition, since, for example, $\mathbb{Z}$ itself has no $\leqslant$-minimal element. (Note, however, that the integers can be well-ordered by, for example, constructing your favorite bijection of sets $h:\mathbb{N}\to\mathbb{Z}$ and ordering the integers by the induced ordering from $h$ and the standard ordering $\leqslant$ on $\mathbb{N}$.) Similarly, the standard ordering $\leqslant$ on the set $0,\infty)$ of non-negative real numbers—which itself has a $\leqslant$-minimal element—fails to satisfy the last condition since, for example, the subset $(0,\infty)$ of positive real numbers has no $\leqslant$-minimal element. (Unfortunately, one must work harder to prove that there is a well-ordering on $0,\infty)$. Indeed, the initial motivation behind Cantor’s well-ordering principle came from the question of whether the set $\mathbb{R}$ of real numbers can be well-ordered—discussed in detail in [Moo82].), without proof, as a valid law of thought. Since his so-called law was not immediately accepted, Cantor found himself seeking a proof. It was in order to provide such a proof that $\mathsf{AC}$ was first explicitly formulated by Ernst Zermelo in 1904 [Zer04].
@@
It is also worth explicitly noting that, even if an argument (in this case, an existence proof) seems to rely on making arbitrary “choices,” $\mathsf{AC}$ may not be needed. For instance, a standard proof of the Bolzano–Weierstrass theorem—every bounded sequence $(a_{i})_{i\in\mathbb{N}}$ in $\mathbb{R}$ contains a convergent subsequence—relies on iteratively “choosing” infinitely many terms $(a_{i_{j}})_{j\in J\subseteq\mathbb{N}}$ from $(a_{i})_{i\in\mathbb{N}}$ satisfying some properties (see [Abb15, Theorem 2.5.5]). In this case, since the indexing set $\mathbb{N}$ is well-ordered by $\leqslant$ (footnote 1), each “choice” can be made definably by choosing the lowest-index term of the sequence satisfying those properties.
@@
The footnotes in this article either define terminology or contain additional information that is not essential to the flow of the paper. A reader who is familiar with the relevant terminology can ignore them and still understand the paper.
We use the abbreviation $\mathsf{ZFC}$ for Zermelo–Fraenkel set theory with $\mathsf{AC}$—which is the standard axiomatic basis for mathematics—and the abbreviation $\mathsf{ZF}$ for Zermelo–Fraenkel set theory without assuming $\mathsf{AC}$. However, familiarity with the axioms of $\mathsf{ZF}$ is not assumed or necessary for this discussion.
@@ 2.2 Mathematical inconveniences
Often, a reason to not believe $\mathsf{AC}$ is its seemingly paradoxical consequences. If there were such a thing as a canonical [^4] See, for example, “the most spectacular [counterintuitive geometrical consequence of $\mathsf{AC}$]” in [Bel09, Ch.I] and “the most stunning [paradoxical result that is demonstrable in $\mathsf{ZFC}$ but not in $\mathsf{ZF}$]” in [Her06, §5.2], as well as [JW96, §9.4: Banach–Tarski paradox] and [Jec73, §1.3: A paradoxical decomposition of the sphere]. example of a “disastrous” consequence of the axiom of choice, it would be the Banach–Tarski paradox—which, roughly stated, says that a closed three-dimensional ball $B$ can be decomposed into finitely many pieces [^5] In fact, five pieces suffice [Rob47]; see [TW16, Ch. 5] or [Hal25, Ch. 7]., which can be rearranged into two disjoint copies of $B$ using only rigid transformations.
@@
Banach–Tarski paradox ($\mathsf{BT}$). Assuming $\mathsf{AC}$, given a closed three-dimensional ball $B$, there exists a decomposition $B=X\sqcup Y$ such that $X~\approx~B~\approx~Y$ [BT24]. [^6] For a detailed discussion on $\mathsf{BT}$, see [TW16].
@@
$\mathsf{BT}$ (hence also $\mathsf{AC}$) implies another mathematically inconvenient result: the existence of subsets of $\mathbb{R}^{n}$, for $n\geqslant 3$, that are not (Lebesgue) measurable. [^7] For a formal and self-contained treatment of the Lebesgue measure on $\mathbb{R}^{n}$, which is beyond the scope of this article, see [SS05, Ch.1]. In the case of $\mathbb{R}$, the intuition behind (Lebesgue) measurability comes from the ordinary notion of length: the measure of a point is zero and that of an open interval $(a,b)\subseteq\mathbb{R}$ is its length, denoted m(a,b)=bam(a,b)=b-a. Since any open subset $U$ of $\mathbb{R}$ can be written uniquely as a disjoint union of open intervals [SS05, Theorem 1.3], the measure of $U$ can be defined as the sum of the lengths of these open intervals. Moreover, the Borel sets—defined as the smallest collection of subsets of $\mathbb{R}$ that includes the open sets and is closed under taking complements and countable unions—are measurable [SS05, Ch.1, §3]. (For instance, since singletons are closed, hence Borel, all countable sets are Borel, hence measurable.) A subset $A$ of $\mathbb{R}$ has measure zero if it can be covered by arbitrarily small unions of open intervals: for any $\varepsilon>0$, there are open sets $U_{i}$ with $\sum_{i}m(U_{i})<\varepsilon$ such that $A\subseteq\bigcup_{i}U_{i}$. (For instance, countably infinite sets have measure zero.) More generally, a set $A\subseteq\mathbb{R}$ is measurable if and only if it it differs from a Borel set by a set of measure zero [SS05, Corollary 3.5]. More generally, as a consequence of $\mathsf{BT}$, there cannot exist any finitely additive measure that is defined on all subsets of $\mathbb{R}^{3}$ and invariant under rigid transformations.
@@
The standard example of a non-measurable subset of $\mathbb{R}$ can be defined as follows. For $x,y\in\mathbb{R}$, we write $x\sim_{\mathbb{Q}}y$ to mean that $x-y\in\mathbb{Q}$, where $\mathbb{Q}$ denotes the set of rational numbers. This is an equivalence relation on $\mathbb{R}$. [^8] A binary relation $\sim$ on a set $X$ is said to be an equivalence relation if it is reflexive (footnote 1), symmetric ($\forall x,y\in X$, if $x\sim y$ then $y\sim x$), and transitive (footnote 1). Given an equivalence relation $\sim$ on $X$, the $\sim$-equivalence class of an element $x\in X$ is the set $x]_{\sim}=\{y\in X\mid x\sim y\}$, and the partition of $X$ by $\sim$ is the set $X/{\sim}=\{[x]_{\sim}\mid x\in X\}$ of all $\sim$-equivalence classes. Since each $\sim$-equivalence class is nonempty (as $x\in[x]_{\sim}$), $\mathsf{AC}$ guarantees the existence of a choice function on the partition $X/{\sim}$. We define a choice set on $X/{\sim}$ as the range of such a choice function. Since distinct equivalence classes are disjoint, a choice set on $X/{\sim}$ will contain exactly one element from each equivalence class. A Vitali set $V$ is a subset of $[0,1]$ containing exactly one element from each $\sim_{\mathbb{Q}}$-equivalence class, or simply a choice set on $0,1]/{\sim_{\mathbb{Q}}}$ (see footnote 8), whose existence is guaranteed by $\mathsf{AC}$. A standard argument of the non-measurability of such a $V$, relying on the translation-invariance of measure, can be found in [SS05, Ch.1, §3].
@@
Note, however, that the existence of a non-measurable set is not provable in $\mathsf{ZF}$, i.e., it is consistent with $\mathsf{ZF}$ (assuming the consistency of inaccessible cardinals) that all subsets of $\mathbb{R}$ are measurable [Sol70]. [^9] For a brief discussion on assumptions that are consistent with $\mathsf{ZF}$ and incompatible with $\mathsf{AC}$, in the context of analysis, see [Sch97, §14]. In this regard, it would seem as though the natural way of eliminating $\mathsf{BT}$ would be to insist that all subsets of $\mathbb{R}$ (hence all subsets of $\mathbb{R}^{n}$ [SS05, Proposition 3.6]) are measurable and to suitably weaken $\mathsf{AC}$. However, this assumption leads to another paradox, which we describe briefly, that is just as unsettling as the one being eliminated.
@@ 3.1 Because a lot of math follows from $\mathsf{AC}$ or needs $\mathsf{AC}$
Thus far, we have seen some uses of $\mathsf{AC}$ that are not necessarily intuitive. For instance, in the introduction, we mentioned that $\mathsf{AC}$ was used to prove the well-ordering principle. In fact, this use is necessary, in the sense that $\mathsf{AC}$ is equivalent to the well-ordering principle. $\mathsf{AC}$ is also regularly used throughout mathematics in another equivalent form, namely Zorn’s Lemma131313“Zorn’s Lemma” is the standard name for this principle. A more accurate title would include the names of Kuratowski and Hausdorff, among others. See [Cam78] for details on the history and origins of the various maximal principles equivalent to $\mathsf{AC}$.: if $(P,\preceq)$ is a nonempty partially ordered set in which every nonempty chain $C\subseteq P$ has a $\preceq$-upper bound, then $P$ has a $\preceq$-maximal element. [^14] A binary relation $\preceq$ on $P$ is said to be a partial order if it is reflexive, anti-symmetric, and transitive (see footnote 1). A subset $C$ of a partially ordered set $(P,\preceq)$ is said to be a chain if the restriction of $\preceq$ to $C$ is total (again, see footnote 1). The equivalence of these statements may not seem obvious at all, as reflected by a famous (ironic) quote of Jerry Bona from 1977: “The Axiom of Choice is obviously true; the Well-Ordering principle is obviously false; and who can tell about Zorn’s Lemma?” [Sch97]. The curious reader, who at this point has surely attempted showing that these statements are equivalent, may find a proof in [Jec73, Theorem 2.1].
@@
Assuming $\mathsf{AC}$, these notions agree: $g$ is continuous at $x$ if and only if $g$ is sequentially continuous at $x$. However, the backward direction of this equivalence cannot be proved in $\mathsf{ZF}$, i.e., it is consistent with $\mathsf{ZF}$ that the backwards direction fails ([Jae65]; discussed in [Moo82, Ch.1.2]).
A natural question to raise is whether we need the “full strength” of $\mathsf{AC}$ for this equivalence to hold. This turns out not to be the case: the backwards direction is equivalent to countable choice from sets of reals, i.e., the assertion that any countable family of nonempty subsets of $\mathbb{R}$ admits a choice function, abbreviated $\mathsf{CC}(\mathbb{R})$ [Her06, Theorem 4.54, 8 and 10]. $\mathsf{CC}(\mathbb{R})$ is the restriction of $\mathsf{AC}$ where $\mathcal{F}$ is countable, i.e., $\lvert\mathcal{F}\rvert\leqslant\lvert\mathbb{N}\rvert$, and each element of $\mathcal{F}$ is a nonempty subset of $\mathbb{R}$. This is an obvious consequence of $\mathsf{AC}$ restricted to countable families of arbitrary nonempty sets, abbreviated $\mathsf{CC}$, which itself is strictly weaker than $\mathsf{AC}$. [^15] Both of the implications $\mathsf{AC}\implies\mathsf{CC}\implies\mathsf{CC}(\mathbb{R})$ are strict. See [Jec73, Ch.8] for the irreversibility of the first implication. Since $\mathsf{CC}(\mathbb{R})$ holds in every permutation model and is a boundable statement (see [HR98, III.§2] and [Jec73, Ch.6])—hence, transferrable to a $\mathsf{ZF}$ model—the irreversibility of the second implication can be seen through any permutation model where $\mathsf{CC}$ fails, such as the Basic Fraenkel Model (see [HR98, III.§2.1] or [Jec73, Ch.4.3]).
Note, furthermore, that $\mathsf{CC}(\mathbb{R})$ implies that the Lebesgue measure (see footnote 7) is countably additive, i.e., if $(A_{i})_{i\in\mathbb{N}}$ are pairwise disjoint measurable subsets of $\mathbb{R}$, then $m(\bigcup_{i}A_{i})=\sum_{i}m(A_{i})$ [Her06, §5.1(E13)]. Countable additivity is not provable in $\mathsf{ZF}$, i.e., it is consistent with $\mathsf{ZF}$ that the Lebesgue measure is not countably additive. [^16] For example, in the Feferman–Lévy Model ([FL63]; see [HR98, III.§1.9]), $\mathbb{R}$ is a countable union of countable sets, so there is no nontrivial countably additive measure on $\mathbb{R}$. (Of course, this in turn means that $\mathsf{CC}(\mathbb{R})$ fails in this model, and so does the aforementioned equivalence of continuity notions.)
@@
Let $F$ be a field and $V$ be an $F$-vector space. If $V$ has a finite spanning set, i.e., if there are $x_{1},\ldots,x_{k}\in V$ such that any $y\in V$ can be written as a linear combination $y=\sum_{i=1}^{k}c_{i}x_{i}$, with $c_{i}\in F$, then it is not hard to show, without appealing to $\mathsf{AC}$, that $V$ has a (finite) basis. Furthermore, if an $F$-vector space $V$ has a finite basis, it is once again not hard to show, without appealing to $\mathsf{AC}$, that any two bases of $V$ must have the same cardinality—so the dimension of $V$ can be defined as the cardinality of any basis of $V$.
@@
For results and discussions on results in other fields of mathematics that rely on $\mathsf{AC}$ or its consequences, we refer the interested reader to [Jec73, Ch.10], [Hod74], [Moo82], [GK91], [HR98] and references there, [Her06, §4], and [HHL+19].